OpenStudy (anonymous):

F(x)= -4x+7 and g(x)= 10x-6 find f(g(x))

OpenStudy (marigirl):

Think of x as 'unit' So lets rewrite. F(x)= -4x+7 which is now -4*'unit' +7 g(x)= 10x-6 which is now 10*'unit' -6 find f(g(x)) = f (10x-6) now lets apply f(x) which is -4*'unit'+7 so now that becomes f(g(x)= -4(10x-6)+7

OpenStudy (anonymous):

ok these are my answers to choose from -40x+24 -40x+31 -40x+64 -40x+70

OpenStudy (marigirl):

now lets expand f(g(x)= -4(10x-6)+7 what do we get?

OpenStudy (anonymous):

-40x+24+7?

OpenStudy (marigirl):

yes now add the 24 and 7

OpenStudy (anonymous):

31

OpenStudy (anonymous):

tytyty

OpenStudy (marigirl):

yes

OpenStudy (anonymous):

i understand now

OpenStudy (marigirl):

no worries, hope my 'unit' methods helped u ..

OpenStudy (anonymous):

yup ty